# Regge trajectories for the hidden bottom tetraquarks in the diquark picture

### v1

@@ III.2 $\lambda-$ and $\rho-$trajectories
We estimate the masses of the hidden bottom tetraquarks for the highly excited states and the masses for the more highly excited states can be calculated easily by the Regge trajectory formula. Although the highly excited states will be unstable, the calculated masses are useful. This is because various models or approaches to calculate the masses of tetraquarks often use different parameter values. These models, despite using different parameter values, often give agreeable predictions for the ground state and the lower excited states. The theoretical predictions for the highly excited states are expected to show differences, which will be instructive.
III.3 Discussions
@@
The Regge trajectories take different form and behave differently in various energy regions Chen:2022flh ; Chen:2021kfw . The $\rho$-mode excitations are in the diquark or in the antidiquark. Both the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q})$ are the heavy-light systems; therefore, the $\rho$-trajectories of the hidden bottom tetraquarks behave as $M{\sim}x_{\rho}^{1/2}$, see Eq. (II.5). The $\lambda$-mode excitations are between diquark $(bq)$ and antidiquark $(\bar{b}\bar{q})$, with the tetraquark $(bq)(\bar{b}\bar{q})$ being the heavy-heavy system; hence, the $\lambda$-trajectories of the hidden bottom tetraquarks behave as $M{\sim}x_{\lambda}^{2/3}$. Discussions on other types of tetraquarks can be found in Ref. plan .
@@
According to Eq. (II.5), there are three series of Regge trajectories: two series of $\rho$-trajectories and one series of $\lambda$-trajectories. There are degeneracy in the $\rho$-trajectories, corresponding to the masses of the $\rho$-mode excited states, as discussed in subsection III.2.
The curvature of the Regge trajectories holds significant importance Chen:2018bbr . In potential models, the curvature is related to the dynamic equation and the confining potential. In Ref. Chen:2023web , we suggest that all Regge trajectories for the diquarks, mesons, baryons and tetraquarks are concave downwards in the $(M^{2},\,x)$ $(x=L,\,N_{r})$ planes. For the light systems, the Regge trajectories are also concave when considering the masses of the light constituents. In fact, the trajectories of baryons and tetraquarks discussed in Refs. Chen:2023web ; Chen:2023djq are the $\lambda$-trajectories. In this work, we show that for the hidden bottom tetraquarks, not only the $\lambda$-trajectories but also the $\rho$-trajectories are concave in the $(M^{2},\,x)$ planes.
IV Conclusions
@@
In this work, we obtain the Regge trajectory formulas for the hidden bottom tetraquarks by using the Regge trajectory formula for the heavy-heavy systems and the newly proposed Regge trajectory relation for the heavy-light diquarks. We apply the Regge trajectory formulas to investigate the hidden bottom tetraquarks which are composed of one heavy-light diquark $(bq)$ and one heavy-light antidiquark $(\bar{b}\bar{q}^{\prime})$. The masses of the $\lambda$-mode excited and $\rho$-mode excited tetraquarks with the hidden bottom are estimated, and the $\lambda$-trajectories and $\rho$-trajectories are discussed.
@@
For the hidden bottom tetraquarks, there are three series of the Regge trajectories: two series of $\rho$-trajectories and one series of $\lambda$-trajectories. They exhibit different behaviors. The $\rho$-trajectories behave as $M{\sim}x_{\rho}^{1/2}$ because the $\rho$-mode excitations are in diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$, which are the heavy-light systems. In contrast, the $\lambda$-trajectories behave as $M{\sim}x_{\lambda}^{2/3}$ because the $\lambda$-mode excitations occur between the diquark and the antidiquark, with the the diquark $(bq)$ and antidiquark $(\bar{b}\bar{q}^{\prime})$ forming a heavy-heavy system $(bq)(\bar{b}\bar{q}^{\prime})$. Moreover, not only the $\lambda$-trajectories but also the $\rho$-trajectories of the hidden bottom tetraquarks are concave downwards in the $(M^{2},\,x)$ plane.
@@ Appendix A States of diquarks
In this section, we list the completely antisymmetric states for the diquarks in $\bar{3}_{c}$ Feng:2023txx .
@@ Appendix B States of tetraquarks
Table 7: The states of tetraquark in the diquark picture. The diquark notation is in 6. Here, $q$ and $q^{\prime}$ represent both the light quarks and the heavy quarks. The superscript $1_{c}$ is omitted.
@@
Appendix C Determination of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes

### v4

@@ II.6 Discussions
By employing the Regge trajectory formulas in (II.2), the masses for the more highly excited states can be calculated easily. We argue that the calculation of the masses of highly excited state is instructive although the highly excited states will be unstable. This is because various models or approaches to calculate the masses of tetraquarks often use different parameter values. These models, despite using different parameter values, often give agreeable predictions for the ground state and the lower excited states. The theoretical predictions for the highly excited states are expected to show differences, which will be instructive.
@@
The Regge trajectories take different form and behave differently in various energy regions Chen:2022flh ; Chen:2021kfw . The hidden charm tetraquark case is similar to the hidden bottom tetraquark case. Both the diquark $(Qq)$ and antidiquark $(\bar{Q}\bar{q}^{\prime})$ are the heavy-light systems; therefore, the $\rho$-trajectories of the hidden bottom and charm tetraquarks behave as $M{\sim}x_{\rho}^{1/2}$, see Eq. (II.2). For the $\lambda$-mode, the tetraquark $(Qq)(\bar{Q}\bar{q}^{\prime})$ is the heavy-heavy system; hence, the $\lambda$-trajectories of the hidden bottom and charm tetraquarks behave as $M{\sim}x_{\lambda}^{2/3}$.
@@
According to Eq. (II.2), there are three series of Regge trajectories: two series of $\rho$-trajectories and one series of $\lambda$-trajectories. Correspondingly, there are three series of masses. When $q=q^{\prime}$, there is degeneracy in the $\rho$-trajectories, corresponding to the degeneracy in masses of the $\rho$-excited states.
For the hidden bottom and charm tetraquarks, both the $\lambda$-trajectories and the $\rho$-trajectories are concave downward in the $(M^{2},\,x)$ planes if the confining potential is linear. In the subsection C.4, we show that for all tetraquarks, not only the $\lambda$-trajectories but also the $\rho$-trajectories are concave downward in the $(M^{2},\,x)$ planes when the masses of the light constituents are considered and the confining potential is linear.
III Conclusions
@@
In this work, we obtain the Regge trajectory formulas for the hidden bottom and charm tetraquarks $(Qq)(\bar{Q}\bar{q}^{\prime})$ by using the Regge trajectory formula for the heavy-heavy systems and the newly proposed Regge trajectory relation for the heavy-light diquarks. We apply the obtained Regge trajectory formulas to investigate the tetraquarks with hidden bottom and charm. The masses of the $\lambda$-mode excited and $\rho$-mode excited tetraquarks are estimated. The calculated results are in agreement with other theoretical predictions. Both the $\lambda$-trajectories and $\rho$-trajectories are discussed.
@@
For the hidden bottom and charm tetraquarks, there are three series of the Regge trajectories: two series of $\rho$-trajectories and one series of $\lambda$-trajectories. They exhibit different behaviors. The $\rho$-trajectories behave as $M{\sim}x_{\rho}^{1/2}$ because the $\rho$-mode excitations are in the heavy-light diquark $(Qq)$ and antidiquark $(\bar{Q}\bar{q}^{\prime})$. In contrast, the $\lambda$-trajectories behave as $M{\sim}x_{\lambda}^{2/3}$ because the $\lambda$-mode excitations occur between the diquark and the antidiquark, with the diquark $(Qq)$ and antidiquark $(\bar{Q}\bar{q}^{\prime})$ forming a heavy-heavy system $(Qq)(\bar{Q}\bar{q}^{\prime})$.
Moreover, the behaviors of the Regge trajectories for various tetraquarks are investigated based on the spinless Salpeter equation. It is shown that both the $\lambda$-trajectories and the $\rho$-trajectories for all tetraquarks are concave downward in the $(M^{2},\,x)$ plane when the masses of the light constituents are considered and the confining potential is linear.
Acknowledgments
We are very grateful to the anonymous referees for the valuable comments and suggestions.
@@ Appendix A States of diquarks
In this section, we list the completely antisymmetric states for the diquarks in $\bar{3}_{c}$ Feng:2023txx , see Table IX.
@@ Appendix B States of tetraquarks
Table 10: The states of tetraquark in the diquark picture. The diquark notation is in 9. Here, $q$ and $q^{\prime}$ represent both the light quarks and the heavy quarks. The superscript $1_{c}$ is omitted.
@@
Appendix C Regge trajectory behaviors for various tetraquarks
In this section, the Regge trajectory behaviors for various tetraquarks are investigated.
C.1 Spinless Salpeter equation
$\omega_{i}=\sqrt{m_{i}^{2}+{\bf p}^{2}}=\sqrt{m_{i}^{2}-\Delta}\;\;(i=1,2).$ (6)
C.2 Regge trajectory relations for various systems
For the heavy-heavy systems, $m_{1},m_{2}{\gg}{{|\bf p|}}$, Eq. (5) reduces to
$+V_{d,t}\Psi_{d,t}({{\bf r}}),$
where
$\mu=m_{1}m_{2}/(m_{1}+m_{2}).$ (10)
By employing the Bohr-Sommerfeld quantization approach Brau:2000st and using Eqs. (7) and (9), we obtain the parametrized relation Chen:2022flh ; Chen:2021kfw
$M=m_{R}+\beta_{x}(x+c_{0x})^{2/3}\,\,(x=l,\,n_{r},\,L,\,N_{r}),$ (11)
with
$\beta_{x}=c_{fx}c_{x}c_{c},\quad m_{R}=m_{1}+m_{2}+C^{\prime},$ (12)
where
$c_{x}$ and $c_{c}$ are
$c_{c}=\left(\frac{\sigma^{\prime 2}}{\mu}\right)^{1/3},\quad c_{l,L}=\frac{3}{2},\quad c_{n_{r},N_{r}}=\frac{\left(3\pi\right)^{2/3}}{2}.$ (15)
$c_{fx}$ are equal theoretically to one and are fitted in practice. In Eq. (11), $m_{1}$, $m_{2}$, $c_{x}$ and $\sigma$ are universal for the heavy-heavy systems. $c_{0x}$ vary with different Regge trajectories.
By employing the Bohr-Sommerfeld quantization approach Brau:2000st and using Eq. (16), the parameterized formula can be written as Chen:2022flh ; Chen:2021kfw
$M=m_{R}+\beta_{x}\sqrt{x+c_{0x}}\;(x=l,\,n_{r},\,L,\,N_{r}).$ (17)
$\beta_{x}$ is in Eq. (12), and with
$c_{c}=\sqrt{\sigma^{\prime}},\quad c_{l,L}=2,\quad c_{n_{r},N_{r}}=\sqrt{2\pi}.$ (18)
For the heavy-light systems, the common choice of $m_{R}$ is Selem:2006nd ; Chen:2021kfw ; Jakhad:2023mni ; Chen:2014nyo ; Chen:2017fcs ; Veseli:1996gy ; Jia:2018vwl
$m_{R}=m_{1}.$ (19)
The usual Regge trajectory, Eq. (17) with (19), is obtained in the limit $m_{1}\to\infty$ and $m_{2}\to 0$. There are different ways of including the light constituent’s mass Selem:2006nd ; Nielsen:2018uyn ; Sonnenschein:2018fph ; MartinContreras:2020cyg ; Chen:2023cws ; Chen:2022flh ; Chen:2023ngj ; Chen:2014nyo ; Afonin:2014nya ; Sergeenko:1994ck . Two modified formulas are proposed in Ref. Chen:2023cws , which can universally describe both the heavy-light mesons and the heavy-light diquarks. One is Eq. (17) with $m_{R}$ in (12), where $m_{2}$ is the light constituent’s mass. Another reads
$M=m_{R}+\sqrt{\beta_{x}^{2}(x+c_{0x})+\kappa_{x}m^{3/2}_{2}(x+c_{0x})^{1/4}}$ (20)
if $m_{2}{\ll}M$, where
$m_{R}=m_{1}+C^{\prime},\quad\kappa_{x}=\frac{4}{3}\sqrt{{\pi}\beta_{x}},$ (21)
where $\beta_{x}$ is in (12). Equation (17) with (12) is an extension of Afonin:2014nya
$M=m_{1}+m_{2}+\sqrt{a(n_{r}+{\alpha}l+b)}$ (22)
and the formula Chen:2022flh
$(M-m_{1}-m_{2}-C)^{2}=\alpha_{x}(x+c_{0})^{\gamma}$ (23)
For the light systems ($m_{1},\,m_{2}\to 0$), Eq. (5) simplifies to
$M\Psi_{d,t}({{\bf r}})=\left[2{{|\bf p|}}+V_{d,t}\right]\Psi_{d,t}({{\bf r}}).$ (24)
By employing the Bohr-Sommerfeld quantization approach Brau:2000st , the parametrized formula can be written as Chen:2022flh ; Chen:2021kfw
$M=\beta_{x}\sqrt{x+c_{0x}}\;(x=l,\,n_{r},\,L,\,N_{r}).$ (25)
For the light systems, the parameters read as
$c_{c}=\sqrt{\sigma^{\prime}},\;c_{l,L}=2\sqrt{2},\;c_{n_{r},N_{r}}=2\sqrt{\pi}.$ (26)
According to the results in Refs. Selem:2006nd ; Sonnenschein:2018fph , similar to Eq. (20) Chen:2023cws , we suggest
$M=C^{\prime}+\sqrt{\beta_{x}^{2}(x+c_{0x})+\kappa_{x}\left(m_{1}^{3/2}+m_{2}^{3/2}\right)(x+c_{0x})^{1/4}}.$ (27)
Based on Eq. (22) Afonin:2014nya and (23), we obtain a modified Regge trajectory relation for the light-light systems in Ref. Chen:2023ngj
$M=m_{R}+\beta_{x}\sqrt{x+c_{0x}},$ (28)
where
$c_{c}=\sqrt{\sigma^{\prime}},\;c_{l,L}=2\sqrt{2},\;c_{n_{r},N_{r}}=2\sqrt{\pi}.$ (29)
$m_{R}$ is in Eq. (12) with $m_{1,2}$ being the masses of the light constituents.
Table 11: The coefficients for the heavy-heavy systems (HHS), the heavy-light systems (HLS), and the light-light systems (LLS).
HHS HLS LLS
$\nu$ $2/3$ $1/2$ $1/2$
$c_{c}$ $\left({\sigma^{\prime 2}}/{\mu}\right)^{1/3}$ $\sqrt{\sigma^{\prime}}$ $\sqrt{\sigma^{\prime}}$
$c_{l,\,L}$ $3/2$ $2$ $2\sqrt{2}$
$c_{n_{r},\,N_{r}}$ ${\left(3\pi\right)^{2/3}}/{2}$ $\sqrt{2\pi}$ $2\sqrt{\pi}$
When Eq. (17) with (12) is employed to discuss the heavy-light systems, and Eq. (28) with (12) is used to discuss the light systems, summarizing Eqs. (11), (17) with (12) and (28) with (12), we have a general form of the Regge trajectories Chen:2022flh ; plan
$M=$ $m_{R}+\beta_{x}(x+c_{0x})^{\nu}\,\,(x=l,\,n_{r},\,L,\,N_{r}),$
$m_{R}=$ $m_{1}+m_{2}+C^{\prime},\quad\beta_{x}=c_{fx}c_{x}c_{c},$ (30)
where ${\nu}$, $c_{x}$ and $c_{c}$ are listed in Table 11. $c_{fx}$ are theoretically equal to one and are fitted in practice. $c_{0x}$ vary with different Regge trajectories. Eq. (C.2) can be employed to discuss various systems including the heavy-heavy systems, the heavy-light systems, and the light-light systems: diquarks, mesons, baryons, and tetraquarks Chen:2023djq ; Chen:2023ngj ; Chen:2023web .
It should be noticed that the general form (C.2) is provisional. Because there are different methods to include the masses of the light constituents. To distinguish the better one, more theoretical and experimental data are needed. In addition to this, the parameter values are universal for both the heavy-heavy systems and the heavy-light systems Chen:2023cws ; Feng:2023txx , while the parameter values should change for the light systems Chen:2023ngj to obtain agreeable results.
[…]
Trajectory behavior $M{\sim}x^{\nu}$ Trajectory behavior $M^{2}{\sim}x^{\nu^{\prime}}$
$\rho_{1}$-mode $\rho_{2}$-mode $\lambda$-mode $\rho_{1}$-mode $\rho_{2}$-mode $\lambda$-mode
1 $(q_{1}q^{\prime}_{1})(\bar{q}_{2}\bar{q}^{\prime}_{2})$ $x_{\rho_{1}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{1}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\rho_{2}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{2}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$ $x_{\rho_{1}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{1}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\rho_{2}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{2}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\lambda}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$
2 $(q_{1}q^{\prime}_{1})(\bar{Q}_{2}\bar{q}_{2})$ $x_{\rho_{1}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{1}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\rho_{2}}^{1/2}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$ $x_{\rho_{1}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{1}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\rho_{2}}^{1/2}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$
3 $(q_{1}Q_{1})(\bar{q}_{2}\bar{q}^{\prime}_{2})$ $x_{\rho_{1}}^{1/2}$ $x_{\rho_{2}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{2}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$ $x_{\rho_{1}}^{1/2}$ $x_{\rho_{2}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{2}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$
4 $(q_{1}q^{\prime}_{1})(\bar{Q}_{2}\bar{Q}^{\prime}_{2})$ $x_{\rho_{1}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{1}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\rho_{2}}^{2/3}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$ $x_{\rho_{1}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{1}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\rho_{2}}^{2/3}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$
5 $(Q_{1}Q^{\prime}_{1})(\bar{q}_{2}\bar{q}^{\prime}_{2})$ $x_{\rho_{1}}^{2/3}$ $x_{\rho_{2}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{2}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$ $x_{\rho_{1}}^{2/3}$ $x_{\rho_{2}}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\rho_{2}}^{3/4}{\,{\boldsymbol{?}}}$ $x_{\lambda}^{1/2}{\,{\boldsymbol{?}}}$, $x_{\lambda}^{2/3}{\,{\boldsymbol{?}}}$
6 $(q_{1}Q_{1})(\bar{Q}_{2}\bar{q}_{2})$ $x_{\rho_{1}}^{1/2}$ $x_{\rho_{2}}^{1/2}$ $x_{\lambda}^{2/3}$ $x_{\rho_{1}}^{1/2}$ $x_{\rho_{2}}^{1/2}$ $x_{\lambda}^{2/3}$
7 $(q_{1}Q_{1})(\bar{Q}_{2}\bar{Q}^{\prime}_{2})$ $x_{\rho_{1}}^{1/2}$ $x_{\rho_{2}}^{2/3}$ $x_{\lambda}^{2/3}$ $x_{\rho_{1}}^{1/2}$ $x_{\rho_{2}}^{2/3}$ $x_{\lambda}^{2/3}$
8 $(Q_{1}Q^{\prime}_{1})(\bar{Q}_{2}\bar{q}_{2})$ $x_{\rho_{1}}^{2/3}$ $x_{\rho_{2}}^{1/2}$ $x_{\lambda}^{2/3}$ $x_{\rho_{1}}^{2/3}$ $x_{\rho_{2}}^{1/2}$ $x_{\lambda}^{2/3}$
9 $(Q_{1}Q^{\prime}_{1})(\bar{Q}_{2}\bar{Q}^{\prime}_{2})$ $x_{\rho_{1}}^{2/3}$ $x_{\rho_{2}}^{2/3}$ $x_{\lambda}^{2/3}$ $x_{\rho_{1}}^{2/3}$ $x_{\rho_{2}}^{2/3}$ $x_{\lambda}^{2/3}$
C.3 Behaviors of the Regge trajectories for various tetraquarks
The Regge trajectories behave differently in different energy regions Chen:2022flh ; Chen:2021kfw . The Regge trajectory behaviors for baryons are discussed in Ref. plan . In this subsection, we discuss the Regge trajectory behaviors for various tetraquarks.
Taken as an example, tetraquarks $(cc)(\bar{q}\bar{q}^{\prime})$ are discussed, where $q$ and $q^{\prime}$ are the light quarks, $u$, $d$, or $s$. The diquark $(cc)$ is the doubly heavy diquark, therefore, the diquark Regge trajectories are clear, see Eq. (11),
$M_{\rho_{1}}{\sim}x^{2/3}_{\rho_{1}}\,\;(x_{\rho_{1}}=l_{1},\,n_{r_{1}}),$ (31)
where $M_{\rho_{1}}$ is mass of the diquark $(cc)$. The antidiquark $(\bar{q}\bar{q}^{\prime})$ is the light diquark. Although they have indefinite forms [see Eqs. (27) and (28) with (12)], the light-diquark Regge trajectories have the same behavior,
$M_{\rho_{2}}{\sim}x^{1/2}_{\rho_{2}}\,\;(x_{\rho_{2}}=l_{2},\,n_{r_{2}}),$ (32)
where $M_{\rho_{2}}$ is mass of the antidiquark $(\bar{q}\bar{q}^{\prime})$. Generally speaking, the diquark Regge trajectories are not the same as the $\rho$-trajectories of tetraquarks.
When discussing the $\lambda$-mode, the Regge trajectory relation for the tetraquarks $(cc)(\bar{q}\bar{q}^{\prime})$ is not clear. We will be confronted with two problems. One is when the light antidiquark can be regarded as being heavy, which will affect not only the behaviors of the ${\lambda}$-trajectories but also the behaviors of the $\rho$-trajectories. Another is how to introduce the mass of the light constituent if the antidiquark is light, which will affect the behavior of the $\rho$-trajectories. If the light antidiquark $(\bar{q}\bar{q}^{\prime})$ is regarded as being light, the tetraquarks $(cc)(\bar{q}\bar{q}^{\prime})$ are the heavy-light systems, therefore, $\lambda$-trajectories behave as
$M{\sim}x^{1/2}_{\lambda}\;(x_{\lambda}=L,\,N_{r})$ (33)
no matter whether Eq. (17) with (12) or Eq. (20) with (21) is chosen as the $\lambda$-trajectory formula. When Eq. (20) is chosen as the $\lambda$-trajectory relation, it will give $M{\sim}M_{\rho_{1}},\,M_{\rho_{2}}^{3/2}$ while Eq. (17) with (12) gives $M{\sim}M_{\rho_{1}},\,M_{\rho_{2}}$ due to different ways to include the light constituent’s mass $M_{\rho_{2}}$ [corresponding to $m_{2}$ in Eq. (17) with (12) and Eq. (20) with (21)]. They give two different behaviors of the $\rho$-trajectories. Using Eqs. (31) and (32), Eq. (17) with (12) gives $M{\sim}x_{\rho_{1}}^{2/3},\,x_{\rho_{2}}^{3/4}$ while Eq. (20) with (21) gives $M{\sim}x_{\rho_{1}}^{2/3},\,x_{\rho_{2}}^{1/2}$. [When discussing the $\rho$-trajectories, highly excited antidiquarks are involved. The first radially and orbitally excited $(ud)$ are about $1.3$ GeV. The masses of the $1^{1}p_{1}$ states of $(us)$ and $(ss)$ are much heavier, $1.46$ GeV and $1.65$ GeV Chen:2023ngj , respectively. They approximate or even are larger than the mass of the charm quark. Therefore, the light antidiquark can be regarded as being heavy.] If the light antidiquark is regarded as being heavy, $\lambda$-trajectories behave as
$M{\sim}x_{\lambda}^{2/3}$ (34)
according to Eq. (11) with (12). Using Eqs. (11), (12), (31), and (32), we have $M{\sim}M_{\rho_{1}},\,M_{\rho_{2}}{\sim}x_{\rho_{1}}^{2/3},\,x_{\rho_{2}}^{1/2}$.
Other types of tetraquarks can be discussed by similar way. The Regge trajectory behaviors of different types of tetraquarks are listed in Table 12.
When the diquark (antidiquark) is doubly heavy or heavy-light, it is clear how to introduce its large mass, see Eqs. (11) with (12) and Eq. (17) with (12) or Eq. (20) with (21). After choosing the Regge trajectory formulas for the heavy-light diquarks, for example, (17) with (12), the Regge trajectory formulas can be written explicitly, and then the behaviors of the $\lambda$-trajectory and the $\rho$-trajectory are definite. Using Eq. (C.2), the Regge trajectory relations for the tetraquarks $(Q_{1}q_{1})(\bar{Q}_{2}\bar{q}_{2})$ read
where
$m_{R{\lambda}}$ $=M_{\rho_{1}}+M_{\rho_{2}}+C,$
$m_{R\rho_{1}}$ $=m_{Q_{1}}+m_{q_{1}}+C/2,$
$m_{R\rho_{2}}$ $=m_{Q_{2}}+m_{q_{2}}+C/2,$
$\mu_{\lambda}$ $=\frac{M_{\rho_{1}}M_{\rho_{2}}}{M_{\rho_{1}}+M_{\rho_{2}}},$
$\beta_{l_{1}}$ $=\sqrt{2\sigma}c_{fl_{1}},\quad\beta_{n_{r_{1}}}=\sqrt{\pi\sigma}c_{fn_{r_{1}}},$
$\beta_{l_{2}}$ $=\sqrt{2\sigma}c_{fl_{2}},\quad\beta_{n_{r_{2}}}=\sqrt{\pi\sigma}c_{fn_{r_{2}}}.$ (36)
The Regge trajectory relations for the tetraquarks $(Q_{1}Q^{\prime}_{1})(\bar{Q}_{2}\bar{q}_{2})$ read
$M_{\rho_{1}}$ $=m_{R\rho_{1}}+\beta_{x_{\rho_{1}}}(x_{\rho_{1}}+c_{0x_{\rho_{1}}})^{2/3}\;(x_{\rho_{1}}=l_{1},\,n_{r_{1}}),$
where
$m_{R{\lambda}}$ $=M_{\rho_{1}}+M_{\rho_{2}}+C,$
$m_{R\rho_{1}}$ $=m_{Q_{1}}+m_{Q^{\prime}_{1}}+C/2,$
$m_{R\rho_{2}}$ $=m_{Q_{2}}+m_{q_{2}}+C/2,$
$\beta_{L}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda}}\right)^{1/3}c_{fL},\;\beta_{N_{r}}=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda}}\right)^{1/3}c_{fN_{r}},$
$\mu_{\lambda}$ $=\frac{M_{\rho_{1}}M_{\rho_{2}}}{M_{\rho_{1}}+M_{\rho_{2}}},\;\mu_{\rho_{1}}=\frac{m_{Q_{1}}m_{Q^{\prime}_{1}}}{m_{Q_{1}}+m_{Q^{\prime}_{1}}},$
$\beta_{l_{1}}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{1}}}\right)^{1/3}c_{fl_{1}},\;\beta_{l_{2}}=\sqrt{2\sigma}c_{fl_{2}},$
$\beta_{n_{r_{1}}}$ $=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{1}}}\right)^{1/3}c_{fn_{r_{1}}},\quad\beta_{n_{r_{2}}}=\sqrt{\pi\sigma}c_{fn_{r_{2}}}.$ (38)
The tetraquark Regge trajectory relations for $(Q_{1}q_{1})(\bar{Q}_{2}\bar{Q}^{\prime}_{2})$ can be obtained easily by the similar way as obtaining the Regge trajectory formulas for $(Q_{1}Q^{\prime}_{1})(\bar{Q}_{2}\bar{q}_{2})$. The Regge trajectory relations for the tetraquarks $(Q_{1}Q^{\prime}_{1})(\bar{Q}_{2}\bar{Q}^{\prime}_{2})$ read
$M_{\rho_{1}}$ $=m_{R\rho_{1}}+\beta_{x_{\rho_{1}}}(x_{\rho_{1}}+c_{0x_{\rho_{1}}})^{2/3}\;(x_{\rho_{1}}=l_{1},\,n_{r_{1}}),$
$M_{\rho_{2}}$ $=m_{R\rho_{2}}+\beta_{x_{\rho_{2}}}(x_{\rho_{1}}+c_{0x_{\rho_{1}}})^{2/3}\;(x_{\rho_{2}}=l_{2},\,n_{r_{2}}),$ (39)
where
$m_{R{\lambda}}$ $=M_{\rho_{1}}+M_{\rho_{2}}+C,$
$m_{R\rho_{1}}$ $=m_{Q_{1}}+m_{Q^{\prime}_{1}}+C/2,$
$m_{R\rho_{2}}$ $=m_{Q_{2}}+m_{Q^{\prime}_{2}}+C/2,$
$\beta_{L}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda}}\right)^{1/3}c_{fL},\;\beta_{N_{r}}=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{\mu_{\lambda}}\right)^{1/3}c_{fN_{r}},$
$\mu_{\lambda}$ $=\frac{M_{\rho_{1}}M_{\rho_{2}}}{M_{\rho_{1}}+M_{\rho_{2}}},\;\mu_{\rho_{1}}=\frac{m_{Q_{1}}m_{Q^{\prime}_{1}}}{m_{Q_{1}}+m_{Q^{\prime}_{1}}},$
$\mu_{\rho_{2}}$ $=\frac{m_{Q_{2}}m_{Q^{\prime}_{2}}}{m_{Q_{2}}+m_{Q^{\prime}_{2}}},\;\beta_{l_{1}}=\frac{3}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{1}}}\right)^{1/3}c_{fl_{1}},\;$
$\beta_{n_{r_{1}}}$ $=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{1}}}\right)^{1/3}c_{fn_{r_{1}}},$
$\beta_{l_{2}}$ $=\frac{3}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{2}}}\right)^{1/3}c_{fl_{2}},$
$\beta_{n_{r_{2}}}$ $=\frac{(3\pi)^{2/3}}{2}\left(\frac{\sigma^{2}}{4\mu_{\rho_{2}}}\right)^{1/3}c_{fn_{r_{2}}}.$ (40)
Due to ambiguity, the tetraquark Regge trajectory formulas for the tetraquarks containing the light diquark or the light antidiquark are not given.
C.4 Discussions
The form of the Regge trajectories depends not only on energy regions but also on the confining potential Chen:2022flh ; Chen:2021kfw . In this work, the employed confining potential is linear. Moreover, various mixings are not considered when discussing the behaviors of the tetraquark Regge trajectories, see Table 12. When the mixings are considered, the behaviors of the tetraquark Regge trajectories will become complex.
In potential models, the curvature of the Regge trajectories is related to the dynamic equation and the confining potential Chen:2018bbr . In Ref. Chen:2023web , we suggest that the Regge trajectories for the diquarks, mesons, baryons and tetraquarks are concave downward in the $(M^{2},\,x)$ planes. In fact, the baryon trajectories and the tetraquark Regge trajectories discussed in Refs. Chen:2023web ; Chen:2023djq actually are the $\lambda$-trajectories. We can see from Table 12 that both the $\lambda$-trajectories and the $\rho$-trajectories for various tetraquarks are all concave downward in the $(M^{2},\,x)$ plane. For the light systems, the trajectories approximate linear as the masses are neglected and become concave when the masses are considered.
If the diquark quark contents have the same flavors as the antidiquark contents, there are degeneracy in masses. The $\rho_{1}$-mode excited states have the same masses as the $\rho_{2}$-mode excited states.
Appendix D Determination of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ for the $\lambda$-modes